OpenStudy (anonymous):

Find the x – intercept of -4x + y = -3 Answer 0 3/4 -3 -3/4

OpenStudy (anonymous):

Get it in the form y = mx +b

OpenStudy (anonymous):

Try it... I'll help you through it :-)

OpenStudy (anonymous):

do i add -4 to both sides?

OpenStudy (anonymous):

How would you get your 4x on the right side?

OpenStudy (anonymous):

Or.. i could add -3 to both sides?/

OpenStudy (anonymous):

y = 4x -3 mx = 4x b = -3

OpenStudy (anonymous):

OHH , ok. thanks hun .

OpenStudy (anonymous):

http://www.mathsisfun.com/equation_of_line.html

OpenStudy (anonymous):

I've got the "teach a person to fish" concept in my head right now and I think it'll help you a lot more since you'll know how to solve all of these on your own :-)

OpenStudy (anonymous):

So the golden rule: always always always get it in the form y = mx +b m = slope b = y-intercept because you'll have your y-intercept when x= 0.

OpenStudy (anonymous):

ok